0NYtiot2 LAUQS 6. Vector A is 52 + 3ý and vector B is 402 +20 a. Find C B/4- 4A b. Convert C into polar coordinates.
Answer #1

a) Å = sê t3 } 3 = 40 et 20% ¿ - i - - - !( toll +zoy) - +(sê + B) 3 loỉt sŷ 20â - 129 ủ 5 - loi - .

a = - 10 sê - Tŷ tano = -1 o - 8² ton? ( J 10²+72 ) = 35° = 12.21 o = 35+ 180 So in Polar From & Lo = 12.21 L215° 2 - 12:21 _****************************************************************************
Goodluck for exam Comment in case any doubt, will reply for sure..